The sides of an equilateral triangle are given as y+5 , 2x-7 and x+y-9 respectively. Find the lengths of the sides.
Answer by swincher4391(1107)

y+5 = x+y-9
5 = x - 9
x = 14
2(14) - 7 = y+5
21 = y+5
y = 16
16 +5 = 21
2(14) -7 = 21
14+16-9 = 21

All the sides are 21.